Lansing Markets just paid an annual dividend of $.70 per share. The firm recently announced that its dividends are expected to increase by 9 percent for the following 3 years and then increase at a constant annual rate of 2.60 percent. If you want to earn 10.50 percent on your investments, how much should you pay for one share of this stock today?
Answer #1

D1=(0.7*1.09)=0.763

D2=(0.763*1.09)=0.83167

D3=(0.83167*1.09)=0.9065203

Value after year 3=(D3*Growth Rate)/(Required rate-Growth rate)

=(0.9065203*1.026)/(0.105-0.026)

=$11.77328896

Hence current price=Future dividends*Present value of discounting factor(rate%,time period)

=0.763/1.105+0.83167/1.105^2+0.9065203/1.105^3+$11.77328896/1.105^3

=$10.77(Approx).